OpenStudy (mony01):

find all points of r=1-cos(theta), r=1+sin(theta)

Parth (parthkohli):

\[-\cos \theta = + \sin\theta\]\[\Rightarrow\tan\theta=-1\]

OpenStudy (mony01):

thats all the points?

Parth (parthkohli):

Find the value of \(\theta\).

Parth (parthkohli):

where \(\tan\theta = -1\)

OpenStudy (mony01):

is it 3pi/4 and 7pi/4

Parth (parthkohli):

Err, I keep going AFK all the time. But yes, that's right.

OpenStudy (mony01):

thank you